OpenStudy (anonymous):

What is the slope of the line that passes through (-1,-1) and (-3,-3)?

OpenStudy (anonymous):

-3+1/-3+1=1

OpenStudy (sloppycanada):

To find the slope, remember that the equation you want to use is \[\frac{ x_2-x_1 }{ y_2 + y_3 }\]

OpenStudy (anonymous):

no formula is y2-y2/x2-x1

OpenStudy (brrandyn):

\[Slope = \frac{ Y _{2} - Y_{1} }{ X _{2} - X_{1} }\] Substitute variables accordingly. \[\frac{ -3 - (-1) }{ -3- (-1)}=\frac{ -2 }{ -2 }= -1\]